### v1

# SoTA with Less: MCTS-Guided Sample Selection for Data-Efficient Visual Reasoning Self-Improvement

Abstract
In this paper, we present an effective method to enhance visual reasoning with significantly fewer training samples, relying purely on self-improvement with no knowledge distillation. Our key insight is that the difficulty of training data during reinforcement fine-tuning (RFT) is critical. Appropriately challenging samples can substantially boost reasoning capabilities even when the dataset is small. Despite being intuitive, the main challenge remains in accurately quantifying sample difficulty to enable effective data filtering. To this end, we propose a novel way of repurposing Monte Carlo Tree Search (MCTS) to achieve that. Starting from our curated 70k open-source training samples, we introduce an MCTS-based selection method that quantifies sample difficulty based on the number of iterations required by the VLMs to solve each problem. This explicit step-by-step reasoning in MCTS enforces the model to think longer and better identifies samples that are genuinely challenging. We filter and retain 11k samples to perform RFT on Qwen2.5-VL-7B-Instruct, resulting in our final model, ThinkLite-VL. Evaluation results on eight benchmarks show that ThinkLite-VL improves the average performance of Qwen2.5-VL-7B-Instruct by 7%, using only 11k training samples with no knowledge distillation. This significantly outperforms all existing 7B-level reasoning VLMs, and our fairly comparable baselines that use classic selection methods such as accuracy-based filtering. Notably, on MathVista, ThinkLite-VL-7B achieves the SoTA accuracy of 75.1, surpassing Qwen2.5-VL-72B, GPT-4o, and O1. Our code, data, and model are available at https://github.com/si0wang/ThinkLite-VL.
Figure 1: Recent “Reasoning VLMs” studies finetune “Base VLMs” with extra reasoning training data to improve visual reasoning. This paper presents a data-efficient self-improving method for better training reasoning VLMs. (Left) Comparison of VLMs with different parameter sizes on MathVista. Our model ThinkLite-VL-7B achieves the state-of-the-art (SoTA) accuracy of 75.1, surpassing Qwen2.5-VL-72B-Instruct, GPT-4o, O1, and other 7B-level reasoning VLMs. (Right) Comparison of the reasoning training data size used by 7B-level reasoning models. Our model achieves SoTA performance using only 11k data, and without any additional knowledge distillation.
1 Introduction
Leveraging long chain-of-thought reasoning with effective reflection during inference, large language models (LLMs) [24, 34] are capable of solving complex reasoning tasks such as math and coding. Recent studies [16] show that large-scale reinforcement fine-tuning (RFT) is a critical factor in enhancing model’s reasoning performance. Notably, substantial reasoning performance improvements can be achieved solely through reinforcement fine-tuning in the post-training stage, even without the standard supervised fine-tuning (SFT) in post-training.
Figure 2: Performance comparison on 8 visual benchmarks. Our model significantly outperforms Qwen2.5-VL-7b-Instruct and other 7b-level reasoning models.
Despite the notable successes in enhancing LLM reasoning with large-scale RFT, similar progress in vision-language models (VLMs) remains limited, likely due to the mismatch between the text-focused pre-training and the multimodal nature of VLM post-training tasks. Recent attempts [22, 12, 53, 81] have employed knowledge-distillation via supervised fine-tuning before the RFT stage, to encourage more visual reasoning related responses being generated. Despite the performance improvement, the knowledge distillation stage is cumbersome, and inherently prevents base VLMs from self-improving themselves in achieving stronger intelligence.
In this paper, we demonstrate that high-quality, appropriately challenging training data is key factor to enable and self-improve visual reasoning ability. When visual reasoning training data aligns properly with the base VLM’s skill level, large-scale RFT alone can significantly enhance VLM’s reasoning ability without relying on knowledge distillation for format fine-tuning or base capability enhancement. Based on this insight, We introduce a data-efficient training pipeline that results in ThinkLite-VL, a reasoning VLM that achieves SoTA visual reasoning performance with less training samples.
The critical factor to ThinkLite-VL’s success is the strategic selection of training samples with suitable difficulty. To achieve this, we repurpose Monte Carlo tree search (MCTS), a classic inference-time search algorithm, to accurately quantify the sample difficulty. Specifically, MCTS’s explicit tree search enforces sufficient thinking compute in deciding the question difficulty, and provide a tight correlation between the question difficulty and the number of MCTS iterations needed to solve it. Our training pipeline begins with collecting 70k open-source samples from three key domains: mathematical reasoning, natural image understanding, and chart comprehension. We then implement MCTS-guided sample selection by applying the VLM itself to perform iterative reasoning on each of the 70k samples, using the number of iterations required to reach the correct solution as a difficulty measure. This rigorous filtering process results in a set of 11k challenging and high-quality samples tailored specifically for our base model. We then directly perform RFT with these selected samples, avoiding any additional supervised fine-tuning steps.
Using the Qwen2.5-VL-7B-Instruct model as our base, we develop our final model, ThinkLite-VL-7B. We evaluate ThinkLite-VL-7B on eight widely used VLM benchmarks. As shown in Figure 2, after RFT with the filtered 11k high-quality data, ThinkLite-VL-7B significantly improves the average performance of Qwen2.5-VL-7B-Instruct from 59.69 to 63.89. It also surpasses the fairly comparable baseline that RFT with the same amount of unfiltered data, from 60.89 to 63.89. Furthermore, compared with the most recent 7B-level reasoning VLMs, ThinkLite-VL-7B consistently demonstrates substantial performance advantages. Notably, on the MathVista benchmark, ThinkLite-VL-7B achieves a state-of-the-art (SoTA) accuracy of 75.1 as shown in Figure 1, significantly surpassing other 7B-level models, open-sourced larger models, GPT-4o, and O1.
2 Related work
Large language model reasoning
Simulating human-like thinking processes through intermediate reasoning steps has significantly improved the performance of large language models (LLMs) on tasks that require reasoning [24]. One family of methods focuses on explicitly controlling the structure or format of the model’s outputs, such as by applying Chain-of-Thought (CoT) prompting [75] and Self-Consistency [74]. Related lines of work include more elaborate reasoning strategies like Tree of Thoughts [84] or Graph of Thoughts [4]. Additionally, some approaches involve supervised fine-tuning (SFT) on curated datasets with reasoning annotations [50, 85]. Researchers have also explored process reward models (PRMs) that encourage systematic thought processes [33, 64, 68, 27, 94, 45]. Others incorporate search techniques, including Monte Carlo Tree Search (MCTS) or beam search, to refine or verify reasoning paths [77, 78, 5, 15, 19, 70]. Recently, large-scale RL with outcome-based reward functions has been leveraged [16] to elicit powerful reasoning capabilities in LLMs. In this paper, we focus on how to use large-scale RL to enhance the reasoning ability of VLMs.
Vision language model reasoning
Vision language models [1, 66, 36, 23, 35, 3, 10, 62, 29, 82] can perform vision tasks using language given visual input through vision encoders like [55, 90, 63]. These models demonstrate comprehensive multimodal capabilities across various scenarios [89, 38, 87, 47, 18, 88, 20, 30] and exhibit reasoning capabilities to some extent [41, 73, 39, 92, 67]. Inspired by the success of reasoning in LLMs, researchers have sought to improve the reasoning capabilities of VLMs. For instance, CoT prompting is applied to VLMs [93, 49, 44, 11, 96, 21] and some papers create multimodal datasets [83, 80, 59, 95, 12, 22, 17, 61], using SFT for knowledge distillation to improve reasoning abilities. Some prior works have also explored improving VLM performance through self-improvement strategies [98, 69, 72, 13]. More recently, RL training has emerged as a promising approach to further strengthen the reasoning capabilities of VLMs [12, 22, 48, 79]. While recent works explore SFT and RL  [12, 22] for VLM reasoning, efficiently utilizing training data and avoiding costly knowledge distillation remains a challenge. In this paper, we propose a novel approach using MCTS to filter for high-quality training instances based on the difficulty level. We then directly apply RL training to enhance reasoning on this curated data, demonstrating strong performance without requiring any SFT stage.
Data filtration
Data filtration aims to identify and retain high-quality, diverse, and task-relevant data while discarding noisy or redundant information to optimize training efficiency and generalization performance. It is important for the pretraining phase [14, 28, 76, 56, 52, 2, 91, 65, 54] and instruction tuning phase [32, 31, 7, 9, 36, 99, 86] of both LLMs and VLMs. In this paper, we specifically focus on filtering training instances to curate data optimally for efficient downstream RL training to improve the reasoning capabilities of VLMs. A concurrent work, MM-Eureka [48], also investigates the impact of data filtration on RFT. However, their approach is limited to a relatively simple self-consistency-based difficulty filtering strategy, where all samples with zero accuracy are discarded. In contrast, we propose a more principled method—MCTS-based sample selection—which enables the identification of truly challenging examples for the VLM. Importantly, our findings reveal that the unsolved samples, which VLMs fail to solve during MCTS, play a critical role in enhancing reasoning performance during RFT, rather than being excluded from the training process.
3 Training Recipe
In this section, we will introduce the complete training pipeline of ThinkLite-VL. First, in Section 3.1, we describe how we collect our training data that we later sample hard problems from. Then, in Section 3.2, we detail how we employ a base model combined with Monte Carlo Tree Search (MCTS) for data filtering to select prompts that are challenging for the base model. Finally, in Section 3.3, we explain how we use these filtered data to train ThinkLite-VL. We note that the proposed data filtering method, introduced in Section 3.2, is the core technical contribution of ThinkLite-VL. Specifically, ThinkLite-VL highlights the importance of difficulty-aware training sample selection in self-improving training, and effectively repurposes MCTS for sample difficulty prediction.
3.1 Data Collection
Category QA Category Data source Data size
Math Reasoning Open-ended Geometry3K 3001
Multi-choice GeoQA 5010
Multi-choice Geos 66
Natural Image Understanding Open-ended FigureQA 10000
Multi-choice ScienceQA 10332
Open-ended OK-VQA 9009
Chart Understanding Open-ended IconQA 10000
Open-ended TabMWP 22579
Figure 3: Data statistic of ThinkLite-VL-70k training dataset. We find that converting all answers to open-ended format is critical in reliably assessing question difficulty and effective model training.
We collect a total of 70k datas from widely used open-source training datasets as our initial training set, covering three category: multimodel mathematical reasoning (Geometry3K [40], GeoQA [6], Geos [58]), natural image understanding (FigureQA [25], ScienceQA [41], OK-VQA [46]), and chart understanding (IconQA [43], TabMWP [42]). For FigureQA and IconQA, due to the large size of their original training sets, we only randomly sample 10k data points from each as our training set. The overall data distribution is shown in Figure 3. Each training sample is organized into the following format: (Image, id, Prompt, Answer).
Furthermore, to prevent the VLM from obtaining correct answers by merely guessing from multiple-choice options, we reformulated IconQA, FigureQA, Geometry3K, TabMWP, and OK-VQA from a multiple-choice format to an open-ended format. This modification compels the VLM to derive the correct answer through reasoning rather than selection, thereby increasing the difficulty of the tasks and enhancing the reliability of the data filtering process described in the subsequent section.
3.2 MCTS-based Sample Selection
Figure 4: Data difficulty distribution of our 11k training set after MCTS-based data filtration. Unsolved refers to data that VLM cannot solve after 50 MCTS iterations.
In our work, the collected data primarily originates from commonly used pretraining datasets for existing VLMs, which makes the model susceptible to overfitting on certain samples. Inspired by recent successes of data filtration in LLM SFT [51, 85] and conventional reinforcement learning [57, 71], we propose a MCTS-based sample selection mechanism. This approach leverages the VLM’s own iterative reasoning process, using the number of iterations required to reach the correct answer as a metric to assess the difficulty of each data sample. Consequently, we can selectively filter for those samples that are more challenging for the model during RL training, rather than using the entire dataset.
Specifically, we define the state at step $t$, denoted as $s_{t}$, to represent the prefix of the reasoning chain. The introduction of a new reasoning step, $a$, transitions the state to $s_{t+1}$, which is formed by concatenating $s_{t}$ with $a$. By leveraging VLM itself as policy model, $\pi_{\theta}$, we sample candidate steps from the probability distribution $\pi_{\theta}(a|x,I,s_{t})$, where $x$ denotes the task’s input prompt and $I$ represents the input image. The MCTS process starts from the root node, $s_{0}$, representing the beginning of a sentence. It then iteratively proceeds through three key phases—selection, expansion and simulation—which are described in detail in the subsequent paragraphs. In contrast to previous studies, during the data filtering stage with MCTS, we prioritize computational efficiency and comprehensive exploration of the solution space, with our focus centered on self-rewarding setting. Consequently, throughout the MCTS process, we do not employ any pretrained or separately trained process reward models, thereby simplifying and accelerating the procedure. The prompt used for MCTS is shown in Table 1.
Table 1: Prompt used for VLM during MCTS procedure. We provide two examples of in-context learning to force VLM to follow the reasoning format.
[…]
Selection
In our MCTS procedure, the selection process is only determined by the visitation frequency, denoted as $N(s_{t})$, of the current state $s_{t}$. At node $s_{t}$, the subsequent node is selected according to the following formula: $s_{t+1}=\arg\max_{s_{t}}\left[\text{c}_{\text{puct}}\cdot\frac{\sqrt{N(s_{t})}}{1+N(s_{t+1})}\right]$
Expansion
Given a current step $s_{t}$, the VLM generates $k$ distinct actions based on the prompt and image through temperature decoding. Each of these actions is then combined with the current step to form $k$ candidates next steps. The diversity among these actions is regulated by temperature parameter, which is set to 0.5 in our experiments, with $k$ configured as 3.
Simulation
After selecting a node , we directly utilize the policy $\pi_{\theta}$ to generate several reasoning steps until a final answer is produced or a preset reasoning step limit is reached. Subsequently, we employ the corresponding LLM (in our experiments, the Qwen2.5-VL-7B-Instruct is used, with Qwen2.5-7B-Instruct serving as the critic model) to compare the generated final answer with the ground truth answer, thereby determining the correctness of the response. If the answer is correct, the MCTS process is terminated and the current iteration number $K$ is recorded; if the answer is incorrect, the visit count $N$ of the selected node is updated and the next iteration commences. Table 2 illustrates the prompt employed for the critic model.
Data filtration
We apply this MCTS procedure to the entire collection of 70k data samples and record the iteration number $K$ required to solve each problem, using Qwen2.5-VL-7B-Instruct as the policy model. In this process, $K$ served as a metric for assessing the difficulty of each sample: a higher $K$ indicates that the VLM requires more extensive exploration to arrive at the correct answer, thereby reflecting a greater level of challenge. Ultimately, we select all samples with $K$ greater than 5, as well as those that remained unsolved after 50 iterations, resulting in a final training set of 11k samples. The data difficulty distribution of this final training set is shown in Figure 4.
Table 2: Critic prompt for MCTS simulation results evaluation.
Critic Prompt Template: Please help me judge the correctness of the generated answer and the corresponding rationale. Question: {} Ground truth answer: {} Generated rationale and answer: {} Your output should only be one sentence: the generated answer is true or false.
3.3 Visual Reasoning Training
Table 3: Visual reasoning training data comparison between ThinkLite-VL and other VLM reasoning models. ALL these reasoning models have distilled knowledge from larger models or closed-source models except for MM-Eureka-Qwen-7B. MM-Eureka-Qwen-7B uses more K12 data (54k) than ours and performs accuracy-based data filtering before training. Here the data size refers to the amount of additional visual reasoning data used to boost the base model for reasoning via SFT or RL training.
Reasoning Models Knowledge Distillation (KD) RFT Data size
LLaVA-Cot-11B [80] GPT-4o 100k
Mulberry-7B [83] GPT-4o, Qwen2-VL-72B 260k
Vision-R1-7B [22] Deepseek-R1 200k + 10k
OpenVLThinker-7B [12] DeepSeek-R1-Distill-Qwen-14B 59.2k
MM-EUREKA-Qwen-7B [48] 54k
ThinkLite-VL-7B 11k
Unlike previous VLM reasoning studies, which heavily depend on large-scale Chain-of-Thought (CoT) data generated by external models and employ SFT for knowledge distillation to enhance reasoning capabilities (as shown in Table 3), we demonstrate that directly performing reinforcement fine-tuning (RFT) with a small amount of high-quality training data can significantly enhance the reasoning ability of VLMs, without the need for extensive external data generation.
After conducting MCTS-based sample selection and obtaining a filtered set of 11k high-quality training data, we then perform RL fine-tuning on the Qwen2.5-VL-7B-Instruct model using these selected data. Specifically, we employ Group Relative Policy Optimization (GRPO) loss function proposed by [60] for training, with the objective defined as follows:
$J_{\text{GRPO}}(\theta)&=\mathbb{E}_{q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi^{\text{old}}_{\theta}(O|q)}\\ &\Bigg{[}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\Bigg{\{}\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi^{\text{old}}_{\theta}(o_{i,t}\mid q,o_{i,<t})}\,\hat{A}_{i,t},\text{clip}\Bigg{(}\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi^{\text{old}}_{\theta}(o_{i,t}\mid q,o_{i,<t})},\,1-\epsilon,\,1+\epsilon\Bigg{)}\,\hat{A}_{i,t}\Bigg{\}}-\beta\,D_{\text{KL}}(\pi_{\theta}\,\|\,\pi_{\text{pre}})\Bigg{]}.\end{aligned}$
We provide the training prompt template during RFT in Table 4.
Table 4: Prompt template used for reinforcement learning fine-tuning.
[…]
4 Experiments
4.1 Benchmark Evaluation
In this subsection, we systematically evaluate ThinkLite-VL on several commonly used multimodal benchmark datasets and perform comprehensive comparisons with existing reasoning models. Through these experiments, we demonstrate the effectiveness and advantages of our model in multimodal reasoning tasks.
Baselines and implementation details
We use Qwen2.5-VL-7B-Instruct as the base model and perform RFT on the 11k high-quality data obtained through MCTS-based filtration, resulting in our proposed model, named ThinkLite-VL-7B. We conduct training using Easy-R1 [97] code base and set GRPO rollout number as 32. Our main baselines are as follows: (1) Qwen2.5-VL-7B-Instruct, serving as our base model; (2) ThinkLite-VL-Random11k, trained using RFT on a randomly sampled subset of 11k instances from the full dataset. Besides, we report the performance of several recent reasoning VLMs for comparison, including the SFT-based models LLaVA-Cot-11B and Mulberry-7B, as well as the RFT-based models Vision-R1, MM-Eureka-Qwen-7B, and OpenVLThinker-7B. We also include larger open-source models and commercial models as SOTA performance references which include Qwen2.5-VL-72B-Instruct, InternVL2.5-78B, GPT-4o, and O1.
Benchmarks
We select eight widely used VLM benchmarks for evaluation, namely MathVista [39], MathVison [67], MathVerse [92], MMMU [89], MMStar [8], MMBench [37], MMVet [87], and AI2D [26]. Among them, MathVista, MathVison, and MathVerse are widely used in VLM research to evaluate mathematical reasoning capabilities, while MMVet also includes a significant number of mathematical reasoning tasks. In contrast, MMMU, MMStar, MMBench, and AI2D are primarily utilized to assess VLM’s visual perception reasoning and scientific reasoning abilities.
Table 5: Comparison of different VLMs on 8 widely used visual benchmarks. The grey sections indicate models with larger parameter sizes and closed-source models. Our model achieves SoTA performance at the 7B level on 6 benchmarks and reaches a SoTA performance of 75.1 on MathVista among all VLMs. On average, our model improves performance by 7% compared with Qwen2.5-VL-7B-Instruct. We do not evaluate Mulberry-7B on MathVision because Mulberry-7B uses MathVision as training dataset, and for Vision-R1-7B, their model is not open-sourced, so we only refer to the results reported in their paper.
Models Data size
MathVista
testmini
MathVision
mini
MathVerse
mini
MMMU
MMStar
MMBench
MM-Vet
AI2D
Avg.
Qwen2.5-VL-72B-Instruct 74.8 39.8 57.6 70.2 70.8 88.6 76.2 88.5 70.81
InterVL2.5-78B 72.3 34.9 51.7 70.1 69.5 88.3 72.3 89.1 68.53
GPT-4o 63.8 36.8 50.2 69.1 64.7 83.4 69.1 84.6 65.21
O1 73.9 78.2
LLaVA-Cot-11B 100k 54.8 16.3 33.9 46.2 57.6 75.0 60.3 78.7 52.85
Mulberry-7B 260k 63.1 39.6 55.0 61.3 79.2 63.7 80.1
Vision-R1-7B 210k 73.5 52.4
OpenVLThinker-7B 59.2k 70.2 29.6 47.9 51.9 63.2 81.3 66.9 82.7 61.71
MM-EUREKA-Qwen-7B 54k 73.0 31.9 50.3 52.3 64.1 79.3 64.9 81.4 62.15
Qwen2.5-VL-7B-Instruct 67.8 23.6 44.5 50.6 61.7 80.7 66.0 82.6 59.69
ThinkLite-VL-Random11k 11k 71.9 26.1 47.3 51.7 62.7 81.1 65.5 80.9 60.89
ThinkLite-VL-7B 11k 75.1 32.9 50.7 54.6 65.0 81.4 67.8 83.6 63.89
SoTA performance over 7B reasoning models
As shown in Table 5, ThinkLite-VL-7B shows a significant improvement in average performance across the eight benchmarks compared to the base model Qwen2.5-VL-7B-Instruct, with the average performance increasing from 59.69 to 63.89. Compared to ThinkLite-VL-Random11k, which is trained with the same data size using random sampling, our method shows significant advantages across all benchmarks, indicating the effectiveness and importance of MCTS-based sample selection. Furthermore, ThinkLite-VL-7B also outperforms reasoning models that primarily achieve performance enhancement through extensive knowledge distillation (such as LLaVA-CoT-11B, Mulberry-7B, Vision-R1-7B, and OpenVLThinker-7B) with the closest average performance to GPT-4o. Compared to MM-EUREKA-Qwen-7B, which does not involve SFT knowledge distillation but adopts a larger RL training dataset, our model consistently outperforms across all benchmarks, highlighting the importance of high-quality data filtering before training, and the effectiveness of the proposed MCTS-based filtering. From the perspective of individual benchmarks, our method achieves the highest scores among 7B-level models on six out of the eight benchmarks. The only exceptions are the MMMU and MathVerse benchmarks, where we slightly lag behind Mulberry-7B and Vision-R1-7B that focused on a narrower range of tasks, respectively. Remarkably, our model achieves the SoTA accuracy of 75.1 on the MathVista benchmark, surpassing larger open-sourced VLMs, GPT-4o, and O1.
4.2 Importance of MCTS-based Sample Selection
In this section, we conduct ablation studies to demonstrate the importance of MCTS-based sample selection. We compare five different training settings of ThinkLite-VL: (1) ThinkLite-VL-Unsolved: Trained using only the 5.6k samples that could not be solved by MCTS, representing the most difficult subset. (2) ThinkLite-VL-Iter5Only: Trained on the subset of data that VLM is able to solve via MCTS, but required more than 5 iterations. This set, combined with the unsolved samples, forms the full 11k training set used in ThinkLite-VL. (3) ThinkLite-VL-Random11k: Trained on a randomly sampled 11k subset from the full 70k dataset, matching the size of the ThinkLite-VL training set. (4) ThinkLite-VL-SelfConsistency: Trained on 23k samples selected based on a self-consistency difficulty measure. Specifically, for each prompt, we perform 50 rollouts using Qwen2.5-VL-7B-Instruct and compute answer accuracy using Qwen2.5-7B-Instruct. Samples with accuracy lower than 0.2 are selected for RFT. (5) ThinkLite-VL-Fullset: Trained on the complete 70k dataset without any filtering. We report the evaluation results of all five settings across the eight VLM benchmarks, as shown in Table 6.
We observe that ThinkLite-VL-7B, trained using 11k samples via MCTS-guided sample selection, achieves the highest average performance (63.89) among all settings. It outperforms not only the random sampling baseline (ThinkLite-VL-Random11k, 60.89) but also models trained on the full dataset (ThinkLite-VL-Fullset, 63.13) and self-consistency-based filtering (ThinkLite-VL-SelfConsistency, 63.15), despite using significantly fewer training samples. This highlights the effectiveness of our difficulty-aware data selection strategy. Further analysis reveals that models trained on subsets derived solely from unsolved samples (ThinkLite-VL-Unsolved, 62.04) or samples requiring more than five iterations (ThinkLite-VL-Iter5Only, 62.38) also show decent performance, suggesting that hard and medium-difficulty samples contribute meaningfully to reasoning ability. However, neither subset alone is sufficient. The combination of both unsolved and medium-difficulty samples yields the strongest and most effective training signal.
Besides, we compare the reward curves during RFT of ThinkLite-VL-Random11k, ThinkLite-VL-Fullset, ThinkLite-VL-Iter5Only, and ThinkLite-VL, as shown in Figure 5. Although ThinkLite-VL-Random11k and ThinkLite-VL-Fullset achieve higher rewards during training, their actual benchmark performances are inferior to ThinkLite-VL. This observation suggests that incorporating a large number of easy samples into training rapidly improves rewards but fails to enhance the model’s reasoning ability. Moreover, ThinkLite-VL exhibits notably lower rewards compared to ThinkLite-VL-Iter5Only, indicating that the unsolved data identified by our MCTS-based sample selection strategy indeed pose significant challenges to the VLM. By progressively learning to solve these challenging problems during training—even if not all are solved completely—the reasoning capabilities of VLMs can be substantially improved.
Table 6: Comparison with models trained on data sampled using different selection strategies, ThinkLite-VL achieves significantly better performance, highlighting the effectiveness and superiority of our proposed MCTS-based sample selection method.
Models Data size
MathVista
testmini
MathVision
mini
MathVerse
mini
MMMU
MMStar
MMBench
MM-Vet
AI2D
Avg.
ThinkLite-VL-7B 11k 75.1 32.9 50.7 54.6 65.0 81.4 67.8 83.6 63.89
ThinkLite-VL-Unsolved 5.6k 73.6 26.9 49.4 52.1 62.7 81.1 67.0 83.5 62.04
ThinkLite-VL-Iter5Only 5.4k 73.5 27.5 50.2 52.5 64.2 80.9 66.9 83.3 62.38
ThinkLite-VL-Random11k 11k 71.9 26.1 47.3 51.7 62.7 81.1 65.5 80.9 60.89
ThinkLite-VL-SelfConsistency 23k 74.6 30.9 50.1 53.8 64.1 81.3 67.1 83.3 63.15
ThinkLite-VL-Fullset 70k 74.3 29.9 52.2 53.1 63.7 81.6 67.2 83.0 63.13
Figure 5: Comparison of reward curves of models trained with different data during RFT. Iter5+Unsolved 11k dataset presents the most challenging learning setting for VLM, highlighting the difficulty of the samples selected by MCTS-based sample selection.
4.3 Ablation Study of Data Difficulty
In this section, we investigate how training data difficulty affects model performance. We present the average performance of models trained using different difficulty data in Table 7. Notably, the model trained with the Iter5+Unsolved subset achieves the highest average score of 63.89, outperforming all other settings. When expanding the difficulty threshold (e.g., Iter10, Iter20, Iter30, and Iter40), the model performance consistently declines, suggesting that medium-difficulty samples are important for improving model reasoning ability. As the difficulty of the training data decreases, the model’s performance also declines. This trend suggests that the inclusion of an excessive number of easy samples may weaken the training signal during RFT and ultimately hurt the model’s reasoning ability.
Table 7: ThinkLite-VL performance under different training data difficulty settings. Iter5+Unsolved achieves the best performance.
Difficulty level Data size Avg. score
Fullset 70k 63.13
Iter1+Unsolved 18k 63.29
Iter5+Unsolved 11k 63.89
Iter10+Unsolved 8k 62.65
Iter20+Unsolved 6.8k 62.61
Iter30+Unsolved 6.1k 62.39
Iter40+Unsolved 5.8k 62.26
Unsolved 5.6k 62.04
5 Case Studies
In this section, we present samples of varying difficulty levels selected by the MCTS-based sample selection method across different datasets, as shown in Tables 13 through 12. The difficulty levels are determined based on the number of reasoning iterations required by the VLM to arrive at the correct answer during the MCTS process, providing reference examples for understanding how the method distinguishes between easy and challenging samples.
Example 3: Different difficulty samples from FigureQA
Iter0 Question: Is Medium Blue less than Dark Orchid?
Ground Truth Answer: Yes.
Iter29 Question: Does Dodger Blue intersect Dark Slate?
Ground Truth Answer: Yes.
Unsolved Question: Does Violet Red have the maximum area under the curve?
Ground Truth Answer: No.
Table 8: Example of samples with different difficulties decided by MCTS-based sample selection from FigureQA.
Example 4: Different difficulty samples from ScienceQA
Iter0 Question: Think about the magnetic force between the magnets in each pair. Which of the following statements is true? Choices: (A) The magnitude of the magnetic force is greater in Pair 2. (B) The magnitude of the magnetic force is greater in Pair 1. (C) The magnitude of the magnetic force is the same in both pairs.
Ground Truth Answer: A.
Iter13 Question: Which solution has a higher concentration of purple particles? Choices: (A) neither; their concentrations are the same (B) Solution A (C) Solution B
Ground Truth Answer: B.
Unsolved Question: What is the direction of this push? Choices: (A) away from the hockey stick (B) toward the hockey stick
Ground Truth Answer: A.
Table 9: Example of samples with different difficulties decided by MCTS-based sample selection from ScienceQA.
Example 5: Different difficulty samples from OK-VQA
Iter0 Question: What food group is pictured here?
Ground Truth Answer: fruit.
Iter20 Question: What is the length of the surfboard the man in the black shorts at the back of the line of people is holding?
Ground Truth Answer: 7 feet.
Unsolved Question: What is this guy’s profession?
Ground Truth Answer: security.
Table 10: Example of samples with different difficulties decided by MCTS-based sample selection from OK-VQA.
Example 6: Different difficulty samples from IconQA
Iter0 Question: How many flowers are there?
Ground Truth Answer: 56.
Iter10 Question: How many dots are there?
Ground Truth Answer: 40.
Unsolved Question: How many stars are there?
Ground Truth Answer: 19.
Table 11: Example of samples with different difficulties decided by MCTS-based sample selection from IconQA.
Example 7: Different difficulty samples from TabMWP
Iter0 Question: Adriana wants to buy 3 pounds of silver confetti. How much will she spend?
Ground Truth Answer: 36.
Iter22 Question: A game show viewer monitors how often a wheel numbered 1 through 5 stops at each number. How many people are there in all?
Ground Truth Answer: 29.
Unsolved Question: The employee at the department store counted the number of ties on each tie rack. How many racks have at least 30 ties but fewer than 70 ties?
Ground Truth Answer: 15.
Table 12: Example of samples with different difficulties decided by MCTS-based sample selection from TabMWP.
Example 1: Different difficulty samples from Geometry3K
Iter0 Question: Find y so that the quadrilateral is a parallelogram.
Ground Truth Answer: 9.
Iter16 Question: Use parallelogram M N P R to find y.
Ground Truth Answer: 6.45.
Unsolved Question: Find the area of the parallelogram. Round to the nearest tenth if necessary.
Ground Truth Answer: 315.
Table 13: Example of samples with different difficulties decided by MCTS-based sample selection from GeoQA.
Example 2: Different difficulty samples from Geos
Iter0 Question: What is the area of the following square, if the length of BD is $2*\sqrt{2}$? Choices: (A) 1 (B) 2 (C) 3 (D) 4 (E) 5.
Ground Truth Answer: D.
Iter7 Question: Given the circle at the right with diameter AB, find x. Choices: (A) 30 degrees (B) 45 degrees (C) 60 degrees (D) 90 degrees (E) None
Ground Truth Answer: D.
Unsolved Question: In the diagram at the right, lines f and g are parallel, and lines a and b are parallel. x = 75. What is the value of y + z? Choices: (A) 75 (B) 105 (C) 150 (D) 180 (E) None
Ground Truth Answer: D.
Table 14: Example of samples with different difficulties decided by MCTS-based sample selection from Geos.
6 Conclusion
We have introduced an effective self-improvement approach to enhance the reasoning capabilities of VLMs, eliminating the need for external supervision or knowledge distillation. Our key insight highlights the critical importance of selecting genuinely challenging examples for Reinforcement Fine-Tuning (RFT). We find that when training data quality is sufficiently high, even a modest dataset can substantially enhance visual reasoning performance without resorting to knowledge distillation methods. Building on this insight, we propose a novel data selection technique, MCTS-based sample selection, which identifies and retains challenging samples by quantifying the number of reasoning iterations required by the VLM to resolve each problem using MCTS. Applying our method to a curated initial set of 70k VLM training samples, we obtain a high-quality subset comprising 11k challenging samples. This curated dataset is then used to fine-tune the Qwen2.5-VL-7B-Instruct model via RFT, resulting in a reasoning VLM named ThinkLite-VL. Our model demonstrates significant improvements across multiple visual reasoning benchmarks, and notably achieves a new SoTA accuracy of 75.1 on MathVista. We hope that our findings on the difficulty-based selection of RFT training data can provide insights for training more effective reasoning VLMs.
Acknowledgment
Wang and Huang are supported by DARPA Transfer from Imprecise and Abstract Models to Autonomous Technologies (TIAMAT) 80321, National Science Foundation NSF-IIS-2147276 FAI, National Science Foundation NAIRR240045, DOD-AFOSR-Air Force Office of Scientific Research under award number FA9550-23-1-0048.